Solve each system b the addition method. Indicate whether it is independent, inconsistent,or dependent.
1/3X-1/6Y=1/3
1/6X+1/4Y=0

Answer by stanbon(75887)

Solve each system b the addition method. Indicate whether it is independent, inconsistent,or dependent.
1st: 1/3X-1/6Y=1/3
2nd: 1/6X+1/4Y= 0
-------------
Multiply thru 1st by 6; Multiply thru 2nd by 12
3rd: 2x - y = 2
4th: 2x + 3y = 0
------------------
Subtract 3rd from 4th to get:
4y=-2
y = -1/2
---------------
Substitute into 2x+3y=0 to solve for x:
2x + 3*(-1/2) = 0
2x = 3/2
x = 3/4
------------------
Final Answer: x=3/4 ; y= -1/2
==========================
Cheers,
Stan H.